UT Sep 04 2022 - Night #3 of Restart



Night Info

  • Sunset 18:41
  • Nautical Twilight Ends 19:36
  • Astro Twilight Ends 20:06
  • Astro Twilight Begins 04:32
  • Nautical Twilight Begins 05:02
  • Sunrise 05:56
  • Moonset 22:40
  • Moon illumination ~55%

Staff

  • OSA: D.G. Huerta
  • ISp: E.Solheid
  • Mtn Manager: B. Waggoner
  • ISA/SO: B.Rothberg, A.Becker, R.T.Gatto
  • Technical Expert: J.Hill

Night Plan

  • Staring Configuration Bino PEPSI Pol
  • LUCI-1 is offline until MCE/ROE powersupply is replaced (or other explanation/fix found)
  • Daytime checks on AGW4

PEPSI Checkout Part Deux

  • Checkout of AGW 3/4 for PEPSI-Pol
    • This is dependent on the results of work on ITs 8701 (AGW 4 issues) and 8702 (Pinhole adjustment for SX) - if not fixed will skip to AGW 1/2 checkout
    • Verify SH spots are well illuminated
    • Send Acquire Preset
    • Active Preset and test collimation
  • PEPI-Pol Flux Calibration Star
  • Name V Inst Fiber BlueArm RedArm Start MST End Duration Phase Moon PID
    gamma Equ 4.70 POL 200P CD3 6x00:04:00 CD5 6x00:04:00 19:20:00 19:48:18 00:28:18 84 polstd
  • HR 7950 3.78 POL 200P CD3 2x00:01:00 CD5 2x00:01:00 19:50:18 19:53:44 00:03:26 71 fluxstd
    • Gamma Equ 21 10 20 +10 07 53 - visible until 02:45 local time
    • HR 7950 20 47 40 -09 29 44 - visible until 01:20 local time
  • Other On-Sky PEPSI/Pol?
  • Reconfigure for PEPSI-PFU - 0.25hr
  • Name V Inst Fiber BlueArm RedArm Start MST End Duration Phase Moon PID
  • HR 7950 3.78 PFU 100 CD3 00:03:00 CD5 00:03:00 19:55:44 20:00:06 00:04:22 70 fluxstd
    HR 7950 3.78 PFU 200 CD3 00:02:00 CD5 00:02:00 20:02:06 20:04:49 00:02:43 70 fluxstd
    HR 7950 3.78 PFU 300 CD3 00:00:40 CD5 00:04:00 20:06:49 20:11:08 00:04:19 70 fluxstd
    • HR 7950 20 47 40 -09 29 44 - visible until 01:20 local time
  • Completed PEPSI-POL SX checkout - PEPSI-POL DX offline

AGW 1 & 2 Checkout Continued

  • verify guiding hotspot center - critical - (1 hr)
    • in /lbt/oacontrol/current/etc have backed up oacontrol.conf to oacontrol.conf.20220904 on oac computer
    • AGW 1
      • Compare with results from previous night:
      • X = 115, Y= 80
      • correct for binning x2 --> X=230, Y=160
      • correct for DETSEC (x+15, y+39) --> X=245, Y=199
      • [lbto@obs1 ~]$ gcsclient getHotspot leftHotspot is (x,y): 249.00, 196.00
      • X offset +4pix, Y offset -3 pix
    • AGW 2

Telescope Wakeup

  • SX M3 Alignment LFBG/DX M3 Alignment RFBG - Critical (3 hrs) - requires LUCI AGWs - both sides can be done in parallel - requires SH spot and hotspot check completed
    • Pupil alignment check - find a star and center it - need Bent Gregorian (1 hr) - need to do before AGW checkout, but active optics needs to be working
    • Pupil wobble (2 hrs) - requires LUCI-1/AGW-1 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform
  • Field Aberration (2 hrs) - Critical (1hour) (SX more important than DX due to SX removal)
    • Can be done in parallel (data collection, analysis in serial)
    • need both AGW1 and AGW2 on the telescope and checked to execute
    • must complete SH spot centering and hotspot check before executing
    • 1.2" or better seeing (1" or better ideal)
    • Target Fields:
      • BS9177 18:39:33.7 +49:05:38.0
      • StoneP1 22:37:45.0 -01:13:40.45
      • BS9188 23:30:33.4 +38:18:57.0
      • BS9107 03:32:03.0 +37:20:40.0

LUCI Wakeup

  • As time permits and if LUCI-1 is avaible (IT 7583 - unable to communicate with MCE/ROE)
    • if time can proceed with LUCI-2 checkout
  • LUCI Scripts Location: Service_Observing/Restart_2022/LUCI
    • LUCI_2022_Restart_V1.74.xml - all scripts (except LS Rot ZP)
    • RotZP _v2.xml - for LS RotZP scripts
  • Reconfigure from PEPSI PFU to LUCIs (0.1 hr)
    • as of 8/31 - imaging (but not PFC) is released
  • Rotator center
  • rotator zeropoint (requires MOS mask) - (0.5 hr)
    • LUCI-1 is empty - use the LS method
    • LUCI-2 has old masks from Feb 2022
      • can use Stone O and Stone P - note that Stone P had issues and required removing 1-2 stars.
    • Masks:
      • ID 903033 - Stone O
      • ID 921934 - Stone A
      • iD 931027 - Stone D
      • Do the LS method Bino to compare L2 with mask and LS method
    • Do a MOS acquisition (imaging of mask and thru-mask image confimration) - verify the theta < 0.1 deg for each mask
    • To update results (if needed)
      • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
    • Longslit Method (backup in case mask exchange(s) aren't completed in time
      • requires target field with two sources in slit, use alignment holes in longslit. Need LMS file with source in longslit (preferably two) - astrometric field ideal
      • use pseudo-LMS files below for acquisition - works in similar manner to MOS acquisition - will need to manually load the LMS files in RTD
  • Collect transform data - (2hrs) - must do RotZP First
    • Target Fields - select one best visible field
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
    • Note the LUCI frame numbers associated with each offset in IDL script
  • Imaging check (1hr)
    • dither in-field
      • NGC 7027 - cycles through all filters (4pt dither) - visible until 4am
    • dither out-of-field (run one)
      • M52 - JHK, dither 5' out of field - visible all night
      • M40 - Broad filters, few narrow filters - visible all night
  • Non-sidereal imaging check (1hr)
    • dither in-field
      • TBD
    • dither out-of-field
      • TBD
  • check longslit - high (1 hr)
    • best to do this after wavelength and PFC cals are completed and released
      • science-like G210 (run one)
        • NGC 6052 -Kband - 1st half
        • NGC 2744 -Kband - 2nd half
      • science-like G200
        • NGC 7027 - zJ and HK - visible until 4am
    • NS longslit (G210 or G200)
      • TBD
  • Photometric Standards - (2hr)
    • min 2 airmasses
    • requires clear skies any seeing - can be done if PFC not released
  • Twilight Flats - low priority


Closed Dome Work


Night Log

Closed Dome

  • UT 00:56 - AGW4 intervention at Horizon did not work. May be able to move something with a hand paddle at zenith - will coordinate with Ilya and Elliott. Arto driving up to mountain. Opening might be delayed.
  • UT 01:02 - polycom discussion with ilya and James R., Elliott will not be going up tonight, but Arto enroute - conditions poor (thick clouds winds) - science will OK delay an opening
  • AllSkyCurrentImage-1.JPG
  • UT 01:44 - thick clouds still overhead (thicker than above image) - closed - hoping for Arto to arrive soon to assist with AGW4 work
  • UT 02:54 - Arto has arrived, will work on AGW4 in the morning (0700hrs) - sill closed for clouds
  • UT 03:09 - reconfiguring for PEPSI-Pol - will do a quick check on AGW3 (acquire, active preset, check SH spots well illuminated, can collimate)
  • ALTA's prediction for tonight's seeing:
  • 20220903_see_evol_time_Height_20_20000.png
  • UT 03:44 opening dome

PEPSI-POL SX ONLY AGW3 Checkout

  • UT 03:51 - moving to pointing star
  • UT 03:59 - collimation preset - seeing ~1.2"
  • UT 04:00 - preset to polarimetric standard star
  • UT 04:30 - exposures complete
  • UT 04:30 - Reconfiguring for bino-LUCIs

LUCIs AGW1 & AGW2 Check Continued

  • AGW rotaor hotspot check
  • UT 04:43 - pointing star check
  • UT 04:48 - error co-pointing limit....thought this was fixed
  • UT 04:49 - still getting co-pointing errors for an Acquire bino preset
  • UT 04:51 - seems to be fixed, NOW sending pointing preset
  • UT04:57 - collimation preset
  • UT 04:59 - starting AGW rotator hotspot check
    • taking 90 sec exposure with gcsclient
    • round circle is very bright and satrated and kind of thick
    • star R=10.93 so also a tad bright
  • UT 05:07 - moving to R=12.1 star to try again
    • seeing on AGW1 1.2" but AGW2 is 1.4" and fainter, did it find the same star?
    • trying this again with AGW1
    • ptincrement 0 CA -10
    • X=112, Y=81
    • x2---> X=224, Y=162
    • correct for DETSEC ---> X=239, Y=201
    • Hotspot is (x,y): 249.00, 196.00
    • offset is X=+10 Y=+5 vs X=+4 Y=-3 yesterday
  • UT 05:15 - closing for high humidity
  • UT 05:25 - humidity doen to 91-93% but the clouds have thickened overhead
  • UT 05:39 - looks like rain on the allsky camera
  • UT 05:50 - allsky camera is now scarily dark....from ground hog day restart to a stephen king atmosphere on the summit....
  • UT 05:54 - fog on the all-sky?
  • UT 05:56 - showers outside - moving to horizon
  • UT 11:37 - B.Rothberg calling it - its past astro twilight and doubt anything will improve before nautical twilight.

-- BarryRothberg - 31 Aug 2022
Topic revision: r8 - 04 Sep 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback